# RLOMM: An Efficient and Robust Online Map Matching Framework with Reinforcement Learning

### v1

@@ Abstract.
Online map matching is a fundamental problem in location-based services, aiming to incrementally match trajectory data step-by-step onto a road network. However, existing methods fail to meet the needs for efficiency, robustness, and accuracy required by large-scale online applications, making this task still a challenging problem. This paper introduces a novel framework that achieves high accuracy and efficient matching while ensuring robustness in handling diverse scenarios. To improve efficiency, we begin by modeling the online map matching problem as an Online Markov Decision Process (OMDP) based on its inherent characteristics. This approach helps efficiently merge historical and real-time data, reducing unnecessary calculations. Next, to enhance the model’s robustness, we design a reinforcement learning method, enabling robust handling of real-time data from dynamically changing environments. In particular, we propose a novel model learning process and a comprehensive reward function, allowing the model to make reasonable current matches from a future-oriented perspective, and to continuously update and optimize during the decision-making process based on feedback. Lastly, to address the heterogeneity between trajectories and roads, we design distinct graph structures, facilitating efficient representation learning through graph and recurrent neural networks. To further align trajectory and road data, we introduce contrastive learning to decrease their distance in the latent space, thereby promoting effective integration of the two. Extensive evaluations on three real-world datasets confirm that our method significantly outperforms existing state-of-the-art solutions in terms of accuracy, efficiency and robustness.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: the ACM International Conference on Management of Data; June 22-27, 2025; Berlin, Germanyisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Since the widespread availability of GPS services in the 1990s, research related to map matching has been a focal point. Map matching is the process of aligning vehicle trajectories onto the actual road network, which is a fundamental problem in location-based services (Kubicka et al., 2018). Online map matching is a sequential process that aligns incoming data incrementally, making it particularly well-suited for real-time applications. This process is a crucial element for various location-based online services and is vital for applications requiring real-time responses such as navigation services (Sud et al., 2008), optimal route planning (Guo et al., 2018; Pedersen et al., 2020), and travel time estimation (Yuan et al., 2022, 2020).
The map matching method can be categorized into two main types: rule-based (Newson and Krumm, 2009; Yang and Gidófalvi, 2018; Wöltche, 2023) and deep-learning-based (Zhao et al., 2019; Feng et al., 2022; Ren et al., 2021; Jiang et al., 2023; Liu et al., 2024). The former treats trajectories as observations and utilizes the Hidden Markov Model (HMM) to infer road segments as hidden states (Eddy, 1996) or uses Markov Decision Process (MDP) with dynamic programming algorithms like value iteration to model offline matching process (Wöltche, 2023). With advancements in deep learning, attention has shifted towards the latter approach, framing the problem as an end-to-end task. This method leverages large datasets to directly learn patterns for matching trajectories to road segments. Overall, both types of methods perform well in terms of matching accuracy in high-sampling-rate offline scenarios. However, when the problem shifts to the online scenario, they exhibit significant inefficiencies due to their inherent nature (Saki and Hagen, 2022; Qu et al., 2023; Chao et al., 2020a; Huang et al., 2021).
@@ 3. Framework
Figure 3. The architecture of our proposed RLOMM, which consists of two parts: online inference and offline training.
@@
2. Encoding Feature: After the construction of the state $s_{i}$, we design the trajectory and road encoding module to obtain effective encodings of the corresponding information within the state. In the trajectory encoding module, we first construct the trajectory transition graph $G_{T}$ based on the set of trajectories, aiming to capture the inter-trajectory correlations to obtain informative trajectory representations. Subsequently, to bridge the interaction between representations of road segments and trajectory, we incorporate the representations of the corresponding road segment for the current trajectory point $\mathcal{T}_{i}$ and apply a graph neural network to derive the trajectory representation. This is then combined with the historical information of trajectory $H_{i}^{t}$ to generate the trajectory encoding $E_{i}^{t}$ for time step $i$. In the design of the road encoding module, we initially employ the road network to construct the link connection graph $G_{R}$. Subsequently, analogous to the trajectory encoding module, we deploy a graph neural network to generate representations for road segments, and in conjunction with the historical information of road $H_{i}^{r}$, we obtain the road encoding $E_{i}^{r}$. Moreover, as mentioned before, the corresponding road segment representation for the current trajectory point $\mathcal{T}_{i}$ is transmitted to the trajectory encoding module. Finally, the hidden states $H_{i+1}^{r}$ and $H_{i+1}^{t}$ produced by the encoding module is retained as the new historical information for the next time step $i+1$.
@@ 4.1. Trajectory & Road Representation
Trajectory Transition Graph Construction. The map matching problem is inherently graph-natured because the road network is represented as a graph composed of nodes (links) and edges (connections). Therefore, representing trajectories in a graphical form as well can unify the representation of the two types of data, thereby alleviating the impact of their heterogeneity. Motivated by this, we further design trajectory transition graph $G_{T}=(V_{T},E_{T})$ to model the trajectories as a graph to address the heterogeneity between roads and trajectories. Specifically, for trajectory set $\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{n}\}$ and trajectory point $g$ mapped to the grid, we construct the node set $V_{T}$ and $E_{T}$, which is defined as follows:
@@
Remark. In terms of the complexity, the trajectory transition graph can be constructed with time and space complexity linear to the number of points in the trajectory set, both of which can be represented as $O(\sum_{i=1}^{n}|\mathcal{T}_{i}|)$, where $|\mathcal{T}_{i}|$ denotes the number of points in $\mathcal{T}_{i}$. This endows the trajectory transition graph with good scalability, enabling our model to be rapidly deployed in practical application scenarios with large amounts of trajectory data, and to incrementally update the graph structure during service provision.
@@ 5.2. Reward Evaluation
Detour Penalty. Unnecessary detours are a significant issue in map matching, often caused by GPS errors where consecutive trajectory points erroneously match upstream of the previous point (Chao et al., 2020b; Shi et al., 2023; Shen et al., 2020; Hu et al., 2023). This issue severely affects the accuracy of matches, leading to abnormal and circuitous results in the matched routes. Motivated by this, the detour penalty is implemented to discourage the model from choosing circuitous or uncharacteristically long paths that deviate significantly from the most direct route between consecutive GPS points.
@@ 5.3. Model Training
Complexity Analysis. We compare RLOMM with MDP-based (Wöltche, 2023), DNN-based (Ren et al., 2021; Jiang et al., 2023; Liu et al., 2024) offline methods, and HMM-based (Newson and Krumm, 2009; Yang and Gidófalvi, 2018; Hu et al., 2023) online method in terms of complexity. Specifically, we analyze the complexity of parameters and computation for the four types of methods, with the results shown in Table 1. Specifically, we model the problem input as follows. The road network contains $n_{R}$ road segments. The number of trajectories is $n_{T}$, the longest trajectory length is $l_{T}$, and each time interval has $a$ new trajectory points. For each trajectory point, the maximum candidates road segments is $c$. The dimension of all hidden representations is $d$.
@@
Dataset Beijing Porto Chengdu
@@ 6.3. Effectiveness of Loss Weight and Partition
Method Beijing Porto Chengdu
@@ 6.5. Efficiency Comparison
Method memory size (MB) training time (s) matching time (s)
@@ 7. Related Work
Map matching problem can be classified to two broad categories (Chao et al., 2020b): Offline Map Matching and Online Map Matching. For Offline Map Matching, researchers focus on matching complete trajectories in offline scenarios, which aims for optimal matching route with less constraint on processing time (Kubicka et al., 2018). For Online Map Matching, trajectory points are continuously sampled and processed in a streaming manner, which imposes high demands on real-time performance.
@@ 7.1. Offline Map Matching
Recently, deep-learning-based map matching methods have attracted the attention of researchers. In order to learn general patterns from trajectory data, (Zhao et al., 2019; Feng et al., 2022) build deep learning based models to utilize all the trajectory data for joint training and knowledge sharing. The authors in (Ren et al., 2021) propose MTrajRec to recover the fine-grained points in trajectories and map match them on the road network in an end-to-end manner. Jiang et al. (Jiang et al., 2023) propose L2MM to generate high-quality representations of low-quality trajectories through high-frequency trajectory enhancement. Liu et al. (Liu et al., 2024) present a graph-based approach that incorporates conditional models to leverage various road and trajectory correlations. As for cellular trajectory map matching, (Shi et al., 2023; Lu et al., 2023; Shen et al., 2024) design different learning methods (e.g., multi-relational graph learning, federated learning, heuristic reinforcement learning optimizer) to learn the mapping function from cellular trajectory points to road segments.
@@ 7.2. Online Map Matching
However, aforementioned online methods are limited by their rule-based nature, which prevents them from fully capturing the spatio-temporal correlations (Qu et al., 2023). Additionally, these methods are predominantly greedy in their matching process, ignoring the impact of current matches on future outcomes, leading to poor robustness and accuracy (Chao et al., 2020a; Huang et al., 2021).

### v2

@@ Abstract.
Online map matching is a fundamental problem in location-based services, aiming to incrementally match trajectory data step-by-step onto a road network. However, existing methods fail to meet the needs for efficiency, robustness, and accuracy required by large-scale online applications, making this task still challenging. This paper introduces a novel framework that achieves high accuracy and efficient matching while ensuring robustness in handling diverse scenarios. To improve efficiency, we begin by modeling the online map matching problem as an Online Markov Decision Process (OMDP) based on its inherent characteristics. This approach helps efficiently merge historical and real-time data, reducing unnecessary calculations. Next, to enhance robustness, we design a reinforcement learning method, enabling robust handling of real-time data from dynamically changing environments. In particular, we propose a novel model learning process and a comprehensive reward function, allowing the model to make reasonable current matches from a future-oriented perspective, and to continuously update and optimize during the decision-making process based on feedback. Lastly, to address the heterogeneity between trajectories and roads, we design distinct graph structures, facilitating efficient representation learning through graph and recurrent neural networks. To further align trajectory and road data, we introduce contrastive learning to decrease their distance in the latent space, thereby promoting effective integration of the two. Extensive evaluations on three real-world datasets confirm that our method significantly outperforms existing state-of-the-art solutions in terms of accuracy, efficiency and robustness.
map matching; reinforcement learning; trajectory; road network
copyright: ccjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 3 (SIGMOD)article: 209publicationmonth: 6doi: 10.1145/3725346ccs: Information systems Spatial-temporal systems
@@ 1. Introduction
Since the widespread availability of GPS services in the 1990s, research related to map matching has been a focal point. Map matching is the process of aligning vehicle trajectories onto the actual road network, which is a fundamental problem in location-based services (Kubicka et al., 2018; Yuan and Li, 2021). Online map matching is a sequential process that aligns incoming data incrementally, making it particularly well-suited for real-time applications. This process is a crucial element for various location-based online services and is vital for applications requiring real-time responses such as navigation services (Sud et al., 2008), optimal route planning (Guo et al., 2018; Pedersen et al., 2020), and travel time estimation (Yuan et al., 2022, 2020; Zheng et al., 2025).
The map matching method can be categorized into two main types: rule-based (Newson and Krumm, 2009; Yang and Gidófalvi, 2018; Wöltche, 2023) and deep-learning-based (Zhao et al., 2019; Feng et al., 2022; Ren et al., 2021; Jiang et al., 2023; Liu et al., 2024). The former treats trajectories as observations and utilizes the Hidden Markov Model (HMM) to infer road segments as hidden states (Eddy, 1996) or uses Markov Decision Process (MDP) with dynamic programming algorithms like value iteration to model offline matching process (Wöltche, 2023). With advancements in deep learning, attention has shifted towards the latter approach, framing the problem as an end-to-end task. This method leverages large datasets to directly learn patterns for matching trajectories to road segments. Overall, both types of methods perform well in terms of matching accuracy in high-sampling-rate offline scenarios. However, when the problem shifts to the online scenario, they exhibit significant inefficiencies due to their inherent nature (Saki and Hagen, 2022; Qu et al., 2023; Chao et al., 2020; Huang et al., 2021).
@@ 3. Framework
Figure 3. The architecture of RLOMM, which consists of two parts: Online Inference and Offline Training.
@@
2. Encoding Feature: After the construction of the state $s_{i}$, we design the trajectory and road encoding module to obtain effective encodings of the corresponding information within the state. In the trajectory encoding module, we first construct the trajectory transition graph $G_{T}$ based on the set of trajectories, aiming to capture the inter-trajectory correlations to obtain informative trajectory representations. Subsequently, to bridge the interaction between representations of road segments and trajectory, we incorporate the representations of the corresponding road segment for the current trajectory point $\mathcal{T}_{i}$ and apply a graph neural network to derive the trajectory representation. This is then combined with the historical information of trajectory $H_{i}^{t}$ to generate the trajectory encoding $E_{i}^{t}$ for time step $i$. In the design of the road encoding module, we initially employ the road network to construct the link connection graph $G_{R}$. Subsequently, analogous to the trajectory encoding module, we deploy a graph neural network to generate representations for road segments, and in conjunction with the historical information of road $H_{i}^{r}$, we obtain the road encoding $E_{i}^{r}$. Moreover, as mentioned before, the corresponding road segment representation for the current trajectory point $\mathcal{T}_{i}$ is transmitted to the trajectory encoding module. Finally, the hidden states $H_{i+1}^{r}$ and $H_{i+1}^{t}$ produced by the encoding module are retained as the new historical information for the next time step $i+1$.
@@ 4.1. Trajectory & Road Representation
Trajectory Transition Graph Construction. The map matching problem is inherently graph-natured because the road network is represented as a graph composed of nodes (links) and edges (connections). Therefore, representing trajectories in a graphical form can also unify the representation of the two types of data, thereby alleviating the impact of their heterogeneity. Motivated by this, we further design trajectory transition graph $G_{T}=(V_{T},E_{T})$ to model the trajectories as a graph to address the heterogeneity between roads and trajectories. Specifically, for trajectory set $\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{n}\}$ and trajectory point $g$ mapped to the grid, we construct the node set $V_{T}$ and $E_{T}$, which is defined as follows:
@@
Remark. In terms of complexity, the trajectory transition graph can be constructed with time and space complexity linear to the number of points in the trajectory set, both of which can be represented as $O(\sum_{i=1}^{n}|\mathcal{T}_{i}|)$, where $|\mathcal{T}_{i}|$ denotes the number of points in $\mathcal{T}_{i}$. This endows the trajectory transition graph with good scalability, enabling our model to be rapidly deployed in practical application scenarios with large amounts of trajectory data, and to incrementally update the graph structure during service provision.
@@ 5.2. Reward Evaluation
Detour Penalty. Unnecessary detours are a significant issue in map matching, often caused by GPS errors where consecutive trajectory points erroneously match upstream of the previous point (Chao et al., 2020; Shi et al., 2023; Shen et al., 2020; Hu et al., 2023). This issue severely affects the accuracy of matches, leading to abnormal and circuitous results in the matched routes. Motivated by this, the detour penalty is implemented to discourage the model from choosing circuitous or uncharacteristically long paths that deviate significantly from the most direct route between consecutive GPS points.
@@ 5.3. Model Training
Complexity Analysis. We compare RLOMM with MDP-based (Wöltche, 2023), DNN-based (Ren et al., 2021; Jiang et al., 2023; Liu et al., 2024) offline methods, and HMM-based (Newson and Krumm, 2009; Yang and Gidófalvi, 2018; Hu et al., 2023) online method in terms of complexity. We analyze the complexity of parameters and computation for the four types of methods, with the results shown in Table 1. Specifically, we model the problem input as follows. The road network contains $n_{R}$ road segments. The number of trajectories is $n_{T}$, the longest trajectory length is $l_{T}$, and each time interval has $a$ new trajectory points. For each trajectory point, the maximum candidates road segments is $c$. The dimension of all hidden representations is $d$.
@@
Beijing Porto Chengdu
@@ 6.3. Effectiveness of Loss Weight and Partition
Beijing Porto Chengdu
@@ 6.5. Efficiency Comparison
memory size (MB) training time (s) matching time (s)
@@ 7. Related Work
Map matching problem can be classified to two broad categories (Chao et al., 2020): Offline Map Matching and Online Map Matching. For Offline Map Matching, researchers focus on matching complete trajectories in offline scenarios, which aims for optimal matching route with less constraint on processing time (Kubicka et al., 2018). For Online Map Matching, trajectory points are continuously sampled and processed in a streaming manner, which imposes high demands on real-time performance.
@@ 7.1. Offline Map Matching
Recently, deep-learning-based map matching methods have attracted the attention of researchers. In order to learn general patterns from trajectory data, (Zhao et al., 2019; Feng et al., 2022) build deep learning based models to utilize all the trajectory data for joint training and knowledge sharing. The authors in (Ren et al., 2021) propose MTrajRec to recover the fine-grained points in trajectories and map match them on the road network in an end-to-end manner. Jiang et al. (Jiang et al., 2023) propose L2MM to generate high-quality representations of low-quality trajectories through high-frequency trajectory enhancement. Liu et al. (Liu et al., 2024) present a graph-based approach that incorporates conditional models to leverage various road and trajectory correlations. As for cellular trajectory map matching, (Shi et al., 2023; Lu et al., 2023; Shen et al., 2024) design different learning methods (e.g., multi-relational graph learning, federated learning, heuristic reinforcement learning optimizer) to learn the mapping function from cellular trajectory points to road segments. As remote sensing imagery is innovatively applied to various spatiotemporal data tasks (Jiang et al., 2024; Chen et al., 2024), it is conceivable that the map matching task could similarly leverage diverse types of novel data in the future.
@@ 7.2. Online Map Matching
However, aforementioned online methods are limited by their rule-based nature, which prevents them from fully capturing the spatio-temporal correlations (Qu et al., 2023). Additionally, these methods are predominantly greedy in their matching process, ignoring the impact of current matches on future outcomes, leading to poor robustness and accuracy (Chao et al., 2020; Huang et al., 2021).
@@ 8. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No.62425203, No.62032003, No.U21B2016). Haitao Yuan is the corresponding author of the work.